# Slot-BERT: Self-supervised Object Discovery in Surgical Video

### v2

@@ 1 Introduction
Research suggests that humans learn to perceive the world through object-specific grouping and association (Kahneman et al., 1992; Tenenbaum et al., 2011). This ability constructs representations akin to object files that bind and track features of an object over time thereby enabling more efficient use by downstream cognitive tasks. Inspired by this cognitive process, unsupervised object-centric learning aims to learn explainable and adaptable representations from unlabeled datasets (Burgess et al., 2019; Greff et al., 2019; Locatello et al., 2020; Greff et al., 2020).
While object-centric learning comes in different forms, one particularly effective apporach incorporates the inductive bias of grouping or binding low-level, unstructured perceptual activations into a set of vectors known as slots. Each slot encapsulates a higher-level compositional entity, such as an object. This grouping or binding process emerges from architectural priors combined with self-supervised learning techniques, including attention and auto-encoding through specialized encoder-decoder architectures. Substantial progress has been made in this field for both image (Seitzer et al., 2023; Fan et al., 2024; Mansouri et al., 2023; Jiang et al., 2023; Wu et al., 2023b) and video (Weis et al., 2020; Kipf et al., 2022; Aydemir et al., 2023; Zadaianchuk et al., 2023; Wu et al., 2023a; Biza et al., 2023; Lee et al., 2024; Qian et al., 2023; Bao et al., 2022b, 2023; Singh et al., 2022, 2024) processing.
@@
These issues become especially pronounced in surgical videos, where temporal dynamics can be highly complex. Instruments and tissues in surgical videos often move at different speeds, and the visibility of specific objects can fluctuate significantly over the course of a procedure. Moreover, many current object-centric learning approaches for videos fail to fully exploit the sequential structure of datasets with longer videos, primarily due to their reliance on RNN-based models (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022) that struggle with learning from long-range temporal coherence. Employing slot-attention mechanisms to process entire video sequences, as proposed in (Singh et al., 2024), could help overcome some challenges but may struggle with scalability for longer videos in real-world datasets due to computational limitations.
@@
In this paper we propose an architecture that is both easy to train and excels in long-range bidirectional temporal reasoning to efficiently handle longer video sequences. We introduce (Slot-BERT), an unsupervised object-centric slot attention model that performs temporal reasoning over slots using a bidirectional transformer encoder. The model treats slots learned from video frames as foundational visual concepts, similar to how text embeddings represent words. By adapting the BERT transformer encoder from natural language processing to process encoded video slots (akin to a sequence of word embeddings), Slot-BERT learns to reconstruct video feature maps, providing implicit supervision for object discovery through masked auto-encoding of slots. This approach implements a temporal bidirectional self-attention mechanism, enabling effective video representation learning. Furthermore, by learning to reconstruct masked slots, Slot-BERT can predict future slots, and we can use this functionality for future initialization of slots that improves the accuracy when applied on longer sequences. We also propose a video slot contrastive loss that increases the independence between slots within a video by maximizing orthogonality in the latent vector space. This regularization leads to a more distinct representation of concepts and increases the precision of segmentation maps.
To summarize, our contributions are as follows: (1) We introduce Slot-BERT, a novel object-centric self-supervised representation learning model based on bidirectional temporal reasoning across video frames using a transformer architecture. (2) We introduce slot-contrastive loss, specifically designed for slot attention, to improve orthogonality between slots. (3) We demonstrate superior temporal coherence, bidirectional reasoning, and zero-shot generalization compared to state-of-the-art methods across four surgical video datasets from three different domains: abdominal, cholecystectomy, and thoracic surgery. (4) Our model addresses the limitations of fully parallelized architectures, scales effectively to longer videos, and can run on affordable hardware.
@@ 2.1 Self-supervised object-centric learning
In video processing, association of slots across time requires initialization and SA on video typically uses RNN-like architectures (Kipf et al., 2022; Zadaianchuk et al., 2023; Singh et al., 2022; Elsayed et al., 2022) that support iterative initialization of slots. These approaches however suffer from instability when training on longer sequences and in scenes with a limited temporal coherence. A parallelized approach, such as (Singh et al., 2024), processes the entire video sequence as a batch, which comes at the cost of computational efficiency and makes it less suitable for real-world video applications. Our SA architecture builds on the RNN framework by incorporating a temporal fusion transformer to enhance long-range bidirectional reasoning. This novel design improves temporal coherence and maintains scalability for longer video sequences while addressing the limitations of prior methods.
@@ 2.2 Masked Information Encoding
Masked-out features or tokens have also been applied to weakly supervised settings, where class labels were provided while the objective was to produce attention maps for detection and segmentation (Hou et al., 2018; Lee et al., 2019). Inspired by BERT, our method treats slots encoded from images as analogous to word embeddings, incorporating a transformer to process video slots. The training objective is to recover image features when slots are partially masked out. Following this strategy, the transformer module can perform reasoning across time instead of merely duplicating information from individual slots.
@@ 3.2 Temporal recurrence of object-centric slot encoder
where is the query vector, is the key vector, is the value vector, and is the attention matrix.
@@ 3.3 Temporal slot transformer
: A learnable positional embedding is added to the slots to encode temporal information:
@@
: To promote robust learning, and allow TST to have bidirectional reasoning ability (i.e., can predict previous and next slots) we employ random masking during training. For a given masking ratio , a subset of frames is masked by setting their corresponding slot values to zero:
@@
Our masked transformer implementation follows previous work in language modeling (Devlin et al., 2019). We use 3 layers of transformer encoder, each with multi-head self-attention (=8), feed-forward layers with hidden dimension 4 and position embedding. The transformer computes attention across all input slots and masked slots , thus modeling dependencies and interactions between unmasked slots effectively. It results in assignments of temporally fused slots .
By leveraging temporal positional embeddings and masking, the temporal slot transformer achieves robust temporal alignment and enhances long-range temporal reasoning. In addition, it can also serve as a future slot prediction module.
@@ 3.4 Slot decoder
: A simple MLP broadcast decoder (Watters et al., 2019) has been used in VAE models and previous slot attention research (Seitzer et al., 2023). Each slot among the slots is broadcast to match the number of spatial patches resulting in tokens for each slot. A learnable positional encoding is added to each token. These tokens are then processed independently using a shared MLP to output reconstructed features and associated alpha masks indicating the slot’s attentive region. The final reconstruction is obtained by a weighted sum:
@@
: The recently introduced SlotMixer decoder (Sajjadi et al., 2022) for 3D object-centric learning has a constant overhead in the number of slots, as it only decodes once per output, requiring less computation. It employs an attention-based approach and operates in three key steps: slots allocation, mixing, and rendering (Sajjadi et al., 2022). The allocation step takes as input the slots and outputs an embedding vector using a cross-attention transformer. The mixing step is similar to a single-head attention step, using the embedding as queries and the slots as keys, to form an attention map , and the slot mix is obtained as
@@ 3.5 Slot Contrastive Learning
To encourage diversity among slots and reduce redundancy, we adopt a contrastive learning loss based on cosine similarity, inspired by the SimCLR framework (Chen et al., 2020b). Unlike SimCLR, which computes similarity for positive pairs, our loss enhances dissimilarity between negative pairs (Radford et al., 2021). Let represent the one of slot vectors in for a given frame. For each slot vector , we compute its cosine similarity with all other slots within the same frame:
@@
where is a temperature parameter. This loss function enhances slot diversity by enforcing the orthogonality between different slot vectors within each frame while preserving their magnitude.
@@ 4.1 Dataset
Cholec80: The Cholec80 dataset (Twinanda et al., 2016) contains 80 cholecystectomy surgery videos downsampled to 1 FPS with presence labels for seven surgical tools in each frame. Following the preprocessing steps of a weakly supervised learning study (Liao et al., 2024), we derived a Transient Object Presence (TOP) video subset from Cholec80 resulting in 5,296 clips with 30 frames each. The test set is based on 100 clips from the CholecSeg8K dataset (Hong et al., 2020), which contains 8,000 frames with instrument and anatomy segmentation from 17 Cholec80 videos.
@@ 4.5 Unsupervised segmentation performance
| MICCAI | DINO-Saur(Seitzer et al., 2023) | 38.20 ± 0.90 | 42.90 ± 0.50 | 62.5 ± 1.3 | 48.40 ± 0.70 | 45.8 ± 2.2 | |
| SAVi(Kipf et al., 2022) | 29.40 ± 0.20 | 33.20 ± 0.10 | 81.7 ± 1.0 | 36.60 ± 0.20 | 40.0 ± 0.5 | ||
| STEVE(Singh et al., 2022) | 27.90 ± 0.20 | 31.50 ± 0.10 | 139.9 ± 0.6 | 34.30 ± 0.10 | 17.0 ± 0.4 | ||
| Slot-Diffusion(Wu et al., 2023b) | 37.50 ± 0.10 | 42.20 ± 0.10 | 70.5 ± 0.2 | 46.30 ± 0.00 | 42.0 ± 0.2 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 46.30 ± 0.40 | 50.10 ± 0.40 | 53.9 ± 1.3 | 55.10 ± 0.50 | 60.1 ± 1.9 | ||
| Ours | 48.90 ± 0.20 | 52.80 ± 0.30 | 44.2 ± 0.6 | 58.20 ± 0.20 | 70.7 ± 0.8 | ||
| Cholec | DINO-Saur(Seitzer et al., 2023) | 25.70 ± 1.00 | 25.50 ± 0.50 | 75.9 ± 0.6 | 33.90 ± 0.70 | 29.1 ± 0.5 | |
| SAVi(Kipf et al., 2022) | 18.90 ± 0.20 | 18.10 ± 0.10 | 107.0 ± 0.2 | 23.70 ± 0.10 | 15.3 ± 0.1 | ||
| STEVE(Singh et al., 2022) | 19.50 ± 0.10 | 18.70 ± 0.00 | 108.9 ± 0.2 | 24.20 ± 0.00 | 19.8 ± 0.7 | ||
| Slot-Diffusion(Wu et al., 2023b) | 12.80 ± 0.00 | 15.70 ± 0.00 | 91.4 ± 0.3 | 19.40 ± 0.00 | 18.9 ± 0.2 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 26.10 ± 0.30 | 25.80 ± 0.20 | 74.9 ± 0.3 | 34.20 ± 0.20 | 34.6 ± 1.9 | ||
| Ours | 28.80 ± 0.30 | 27.80 ± 0.40 | 64.8 ± 0.9 | 37.00 ± 0.60 | 35.3 ± 2.0 |
@@
The performance on the Cholec dataset is comparatively lower for all models. Nevertheless, our method outperforms all baselines, with notable gains over Video-Saur (e.g., +2.7% in MBO-V, +2.8% in FG-ARI and -10.1 in mBHD). It is evident that the smaller size of the dataset and less diversity in the data impact performance compared to MICCAI, the latter being almost five times larger than Cholec. An additional challenge in Cholec dataset is the frequent instrument disappearance/reappearance (Liao et al., 2024) and the small size of some objects which pose a challenge because the slot number is fixed. These data characteristics also result in the phenomenon where mBO-F is smaller than mBO-V, unlike in MICCAI data, where most methods achieve a higher mBO-F instead.

### v3

@@ 1 Introduction
Research suggests that humans learn to perceive the world through object-specific grouping and association (Kahneman et al., 1992; Tenenbaum et al., 2011). This ability constructs representations akin to object files that bind and track features of an object over time thereby enabling more efficient use by downstream cognitive tasks. Inspired by this cognitive process, self-supervised object-centric learning aims to learn explainable and adaptable representations from unlabeled datasets (Burgess et al., 2019; Greff et al., 2019; Locatello et al., 2020; Greff et al., 2020).
While object-centric learning comes in different forms, one particularly effective approach incorporates the inductive bias of grouping or binding low-level, unstructured perceptual activations into a set of vectors known as slots. Each slot encapsulates a higher-level compositional entity, such as an object. This grouping or binding process emerges from architectural priors combined with self-supervised learning techniques, including attention and auto-encoding through specialized encoder-decoder architectures. Substantial progress has been made in this field for both image (Seitzer et al., 2023; Fan et al., 2024; Mansouri et al., 2023; Jiang et al., 2023; Wu et al., 2023b) and video (Weis et al., 2020; Kipf et al., 2022; Aydemir et al., 2023; Zadaianchuk et al., 2023; Wu et al., 2023a; Biza et al., 2023; Lee et al., 2024; Qian et al., 2023; Bao et al., 2022b, 2023; Singh et al., 2022, 2024) processing.
@@
These issues become especially pronounced in learning disentangled representation for scene decomposition in surgical videos, where temporal dynamics can be highly complex. Instruments and tissues in surgical videos often move at different speeds, and the visibility of specific objects can fluctuate significantly over the course of a procedure. Moreover, many current object-centric learning approaches for videos fail to fully exploit the sequential structure of longer episodes. In surgical videos, a typical episode depicting a surgical action or a surgical task will last from a few seconds to a minute or longer. RNN-based models (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022) struggle with learning from sequences of this length. Employing slot-attention mechanisms to process entire video sequences in parallel, as proposed in (Singh et al., 2024), could help overcome some challenges but may struggle with scalability for longer videos in real-world datasets due to computational limitations.
These limitations underscore the need for an architecture that is computationally tractable for longer video sequences and supports self-supervised training without auxiliary cues which are not inherently robust (e.g., optical flow or between-frame patch similarity). An ideal model would operate in the latent space that supports scalable temporal reasoning well beyond the short effective context of auto-regressive RNNs.
@@
We propose an architecture that is both easy to train and excels in long-range bidirectional temporal reasoning to efficiently handle longer video sequences — Slot-BERT, an self-supervised object-centric slot attention model that performs temporal reasoning over slots using a bidirectional transformer encoder. The model treats slots learned from video frames as foundational visual concepts, similar to how text embeddings represent words. By adapting the BERT transformer encoder from natural language processing to process encoded video slots (akin to a sequence of word embeddings), Slot-BERT learns to reconstruct video feature maps, providing implicit supervision for object discovery through masked auto-encoding of slots. This approach implements a temporal bidirectional self-attention mechanism, enabling effective video representation learning. Furthermore, by learning to reconstruct masked slots, Slot-BERT can predict future slots, and we can use this functionality for future initialization of slots that improves the accuracy when applied on longer sequences. We also propose a video slot contrastive loss that increases the independence between slots within a video by maximizing orthogonality in the latent vector space. This regularization leads to a more distinct representation of concepts and increases the precision of segmentation maps.
To summarize, our contributions are as follows:
We introduce Slot-BERT, a novel object-centric self-supervised representation learning model based on bidirectional temporal reasoning across video frames.
We introduce slot-contrastive loss, specifically designed for slot attention, to improve orthogonality between slots.
Our model is computational efficient and easy to train on longer videos, and can run on affordable hardware.
We demonstrate superior temporal coherence, bidirectional reasoning, and zero-shot generalization compared to state-of-the-art methods across four surgical video datasets from three different domains.
@@ 2.1 Self-supervised object-centric learning
In video processing, association of slots across time requires initialization and SA on video typically uses RNN-like architectures (Kipf et al., 2022; Zadaianchuk et al., 2023; Singh et al., 2022; Elsayed et al., 2022) that support iterative initialization of slots. These approaches however suffer from instability when training on longer sequences and in scenes with a limited temporal coherence. A parallelized approach, such as (Singh et al., 2024), processes the entire video sequence as a batch, which comes at the cost of computational efficiency and makes it less suitable for real-world video applications.
Our SA architecture builds on the RNN framework by incorporating a temporal fusion transformer to enhance long-range bidirectional reasoning, which is inspired by BERT for language modeling (Devlin et al., 2019). We treat the slot representation as a language token and adopt the same pre-training strategy that is used in training of language models to the training of video SA models. This novel design improves temporal coherence and maintains scalability for longer video sequences while addressing the limitations of prior methods.
@@ 2.2 Masked Information Encoding
Masked-out features or tokens have also been applied to weakly supervised settings, where class labels were provided while the objective was to produce attention maps for detection and segmentation (Hou et al., 2018; Lee et al., 2019).
Inspired by BERT, our method treats slots encoded from images as analogous to word embeddings, incorporating a transformer to process video slots. The training objective is to recover image features when slots are partially masked out. Following this strategy, the transformer module can perform reasoning across time instead of merely duplicating information from individual slots.
@@ 3.2 Temporal recurrence of object-centric slot encoder
where is the query vector, is the key vector, is the value vector, and is the attention matrix. is the normalized attention used for aggregation,, is the slots index, indexes the input features, and is the summation index over features.
@@ 3.3 Temporal slot transformer
A learnable positional embedding is added to the slots to encode temporal information:
@@
To promote robust learning, and allow TST to have bidirectional reasoning ability (i.e., can predict previous and next slots) we employ random masking during training. For a given masking ratio , a subset of frames is masked by setting their corresponding slot values to zero:
@@
Our masked transformer implementation follows previous work in language modeling (Devlin et al., 2019). The key to bi-directionality is that the self-attention mechanism is non-causal, in contrast to RNN-like models. Self-attention computes dependencies across all time steps simultaneously, allowing each slot to attend to both past and future frames. During training, the model is tasked with reconstructing the original from , which encourages learning temporally coherent representations that exploit information from the entire sequence in both directions. This design explicitly enables reasoning in both temporal directions.
We use 3 layers of transformer encoder, each with multi-head self-attention (=8), feed-forward layers with hidden dimension 4 and position embedding. The transformer computes attention across all input slots and masked slots , thus modeling dependencies and interactions between unmasked slots effectively. It results in assignments of temporally fused slots . By leveraging temporal positional embeddings and masking, the temporal slot transformer achieves robust temporal alignment and enhances long-range temporal reasoning. In addition, it can also serve as a future slot prediction module.
@@ 3.4 Slot decoder
A simple MLP broadcast decoder (Watters et al., 2019) has been used in VAE models and previous slot attention research (Seitzer et al., 2023). Each slot among the slots is broadcast to match the number of spatial patches resulting in tokens for each slot. A learnable positional encoding is added to each token. These tokens are then processed independently using a shared MLP to output reconstructed features and associated alpha masks indicating the slot’s attentive region. The final reconstruction is obtained by a weighted sum:
@@
The recently introduced SlotMixer decoder (Sajjadi et al., 2022) for 3D object-centric learning has a constant overhead in the number of slots, as it only decodes once per output, requiring less computation. It employs an attention-based approach and operates in three key steps: slots allocation, mixing, and rendering (Sajjadi et al., 2022). The allocation step takes as input the slots and outputs an embedding vector using a cross-attention transformer. The mixing step is similar to a single-head attention step, using the embedding as queries and the slots as keys, to form an attention map , and the slot mix is obtained as
@@ 3.5 Slot Contrastive Learning
To encourage diversity among slots and reduce redundancy, one possible approach is to adopt clustering-based metrics such as Normalized Mutual Information for designing the loss function. However, inspired by the SimCLR framework (Chen et al., 2020b), we adopt a contrastive learning loss based on cosine similarity , which is more computationally efficient, directly optimizable during training, and guarantees differentiability. Unlike SimCLR, which computes similarity for positive pairs, our loss enhances dissimilarity between negative pairs (Radford et al., 2021). Unlike conventional methods that use contrastive loss to learn a single global image representation, we apply it to encourage dissimilarity among slot representations that decompose the image and represent objects across video frames. Let represent the one of slot vectors in for a given frame. For each slot vector , we compute its cosine similarity with all other slots within the same frame:
@@
where is a temperature parameter. By minimizing this loss the cosine similarity between slots is reduced, which will push the slot vectors to different directions in the representation hyper-space, thus increasing the distance between different slot vectors that now lie on a sphere and point in quasi-orthogonal directions.
@@ 4.1 Dataset
Cholec80: The Cholec80 dataset (Twinanda et al., 2016) contains 80 cholecystectomy surgery videos downsampled to 1 FPS with presence labels for seven surgical tools in each frame. Following the preprocessing steps of a weakly supervised learning study (Liao et al., 2024), we derived a Transient Object Presence (TOP) video subset from Cholec80 resulting in 5,296 clips with 30 frames each, in which objects are not necessarily present in every frame, but can rather transition in or out of the field of view. The test set is based on 100 clips from the CholecSeg8K dataset (Hong et al., 2020), which contains 8,000 frames with instrument and anatomy segmentation from 17 Cholec80 videos.
@@
Among these four datasets, only the derivative of Cholec80, CholecSeg8k provides both instrument and tissue ground truth masks for evaluation, while other datasets provide only instrument masks.
To demonstrate the benefit of applying our method to broader domains, We also evaluate Slot-BERT on non-surgical datasets, including a natural real-world video dataset YT-VIS (Xu et al., 2019) and a synthetic multi-object tracking data MOVi-E (Greff et al., 2022). Results on these two datasets are concluded in Supplementary material F.
@@ 4.5 Unsupervised segmentation performance
| MICCAI | DINO-Saur(Seitzer et al., 2023) | 38.2 ± 0.9 | 42.9 ± 0.5 | 62.5 ± 1.3 | 48.4 ± 0.7 | 45.8 ± 2.2 | |
| SAVi(Kipf et al., 2022) | 29.4 ± 0.2 | 33.2 ± 0.1 | 81.7 ± 1.0 | 36.6 ± 0.2 | 40.0 ± 0.5 | ||
| STEVE(Singh et al., 2022) | 27.9 ± 0.2 | 31.5 ± 0.1 | 139.9 ± 0.6 | 34.3 ± 0.1 | 17.0 ± 0.4 | ||
| Slot-Diffusion(Wu et al., 2023b) | 37.5 ± 0.1 | 42.2 ± 0.1 | 70.5 ± 0.2 | 46.3 ± 0.0 | 42.0 ± 0.2 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 46.3 ± 0.4 | 50.1 ± 0.4 | 53.9 ± 1.3 | 55.1 ± 0.5 | 60.1 ± 1.9 | ||
| Ours | 48.9 ± 0.2 | 52.8 ± 0.3 | 44.2 ± 0.6 | 58.2 ± 0.2 | 70.7 ± 0.8 | ||
| Cholec | DINO-Saur(Seitzer et al., 2023) | 25.7 ± 1.0 | 25.5 ± 0.5 | 75.9 ± 0.6 | 33.9 ± 0.7 | 29.1 ± 0.5 | |
| SAVi(Kipf et al., 2022) | 18.9 ± 0.2 | 18.1 ± 0.1 | 107.0 ± 0.2 | 23.7 ± 0.1 | 15.3 ± 0.1 | ||
| STEVE(Singh et al., 2022) | 19.5 ± 0.1 | 18.7 ± 0.0 | 108.9 ± 0.2 | 24.2 ± 0.0 | 19.8 ± 0.7 | ||
| Slot-Diffusion(Wu et al., 2023b) | 12.8 ± 0.0 | 15.7 ± 0.0 | 91.4 ± 0.3 | 19.4 ± 0.0 | 18.9 ± 0.2 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 26.1 ± 0.3 | 25.8 ± 0.2 | 74.9 ± 0.3 | 34.2 ± 0.2 | 34.6 ± 1.9 | ||
| Ours | 28.8 ± 0.3 | 27.8 ± 0.4 | 64.8 ± 0.9 | 37.0 ± 0.6 | 35.3 ± 2.0 | ||
| Cholec + Tissue Mask | DINO-Saur(Seitzer et al., 2023) | 30.2 ± 0.1 | 36.7 ± 0.5 | 90.9 ± 1.0 | 41.4 ± 0.6 | 22.5 ± 0.2 | |
| SAVi(Kipf et al., 2022) | 27.8 ± 0.0 | 29.3 ± 0.0 | 117.2 ± 0.1 | 35.4 ± 0.1 | 16.4 ± 0.0 | ||
| STEVE(Singh et al., 2022) | 28.5 ± 0.0 | 33.6 ± 0.1 | 127.7 ± 0.0 | 33.0 ± 0.0 | 16.5 ± 0.1 | ||
| Slot-Diffusion(Wu et al., 2023b) | 29.3 ± 0.2 | 34.0 ± 0.2 | 99.1 ± 0.2 | 35.5 ± 0.2 | 23.3 ± 0.2 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 30.9 ± 0.1 | 33.7 ± 0.1 | 90.3 ± 0.1 | 43.1 ± 0.3 | 24.8 ± 0.3 | ||
| Slot-BERT (Ours) | 33.4 ± 0.5 | 40.2 ± 0.3 | 85.4 ± 0.8 | 44.1 ± 0.5 | 25.9 ± 0.3 |
† We report Cholec results on instruments only (Cholec) for easier comparison to MICCAI, and on both instruments and tissue (Cholec + Tissue Masks).
@@
The performance on the Cholec dataset is comparatively lower for all models. Nevertheless, our method outperforms all baselines, with notable gains over Video-Saur (e.g., +2.7% in MBO-V, +2.8% in FG-ARI and -10.1 in mBHD on instrument segmentation; +2.7% in MBO-V when evaluated with tissue masks ). It is evident that the smaller size of the dataset and less diversity in the data impact performance compared to MICCAI, the latter being almost five times larger than Cholec. An additional challenge in Cholec dataset is the frequent instrument disappearance/reappearance (Liao et al., 2024) and the small size of some objects which pose a challenge because the slot number is fixed. These data characteristics also result in the phenomenon where mBO-F is smaller than mBO-V, unlike in MICCAI data, where most methods achieve a higher mBO-F instead. Figure 2 presents qualitative results of unsupervised segmentation, illustrating our method’s advantage in delineating instruments from tissue and discriminating tissue with different textures compared to STEVE and SAVi.